one more question!!

find h when v = 141.3 and r = 3 use pi = 3.14

Answer by checkley71(8403)

ASSUMING THE FORMULA IS V=PI*R^2*H
THEN:
141.3=3.14*3^2*H
H=141.3/3.14*9
H=141.3/28.26
H=5 IS THE HEIGHT.